OpenStudy (ballislife):

How do you find the slope and y-intercept of -5x=8-y

OpenStudy (steve816):

First, you must get it into y=mx+b form. Do you think you can do that?

pooja195 (pooja195):

We need to turn this equation into slope intercept form y=mx+b -5x=8-y What can you do to get y by itself?

OpenStudy (ballislife):

subtract 8 on both sides

pooja195 (pooja195):

Good what do you get once you do that? (what equation)

OpenStudy (ballislife):

divide -1 on both sides

pooja195 (pooja195):

ok so far we have -y=-5x-8 Yep divide by -1

pooja195 (pooja195):

Remeber that we need to divide -1 by EVERYTHING.

OpenStudy (ballislife):

would it be 5x+8=y

pooja195 (pooja195):

Yep But lets keep it in this order so it is easier to identify the slope y=5x+8 y=mx+b m=slope So what is the slope?

OpenStudy (ballislife):

5 and the y-intercept is 8, right?

pooja195 (pooja195):

Perfect :)

OpenStudy (ballislife):

thanks

pooja195 (pooja195):

yw!